# ZeroDVFS: Zero-Shot LLM-Guided Core and Frequency Allocation for Embedded Platforms

### v1

@@ Abstract
times\$ faster than table-based profiling, enabling practical deployment in dynamic embedded systems.
@@ 1 Introduction
Energy-efficient embedded systems are limited by computational resources and play a crucial role in various applications, including IoT devices, wearable electronics, and autonomous vehicles. In these systems, managing the thermal dynamics prevents thermal overheating, ensures system reliability, and avoids non-uniform aging due to uneven workload distribution among the cores [9, 19]. However, the available thermal-aware energy-efficient methods proposed for embedded systems lack scalability to many cores and adaptability to different platforms, especially in dynamic and resource-constrained environments, and often incur prohibitive computational overhead. Complementary approaches in the literature address related challenges through hierarchical multi-agent scheduling, statistical feature-aware task allocation, few-shot reinforcement learning, and graph-driven performance modeling.
@@
Given the complexity of embedded systems, the cost-intensive nature of collecting real-world data adds to the challenges of task and processor profiling. This complexity frequently leads to unpredictable execution time estimates, aggravated by limited hardware feedback mechanisms within these systems. These challenges can significantly hinder efficient core allocation, affecting overall performance and energy efficiency. Consequently, existing task scheduling strategies mostly prioritize workload demand or only utilization metrics over temperature and system-related considerations. This trend is discernible in Linux kernel governors [9] and recently formulated policies [38, 30, 41]. Consequently, these strategies often overlook detailed insights provided by system profilers during task execution [3] and allocate resource-intensive tasks to hot cores, which are more sensitive to temperature elevations. State-of-the-art heuristic approaches like the precise scheduler [7] achieve near-optimal energy efficiency for static frequency assignment through exhaustive offline profiling across all configurations, but require 8 to 12 hours of table generation per benchmark (e.g., profiling a 1 to 5 second task across all core-frequency combinations with multiple repetitions), preventing runtime adaptation. Their static nature limits adaptability to thermal variations and workload phase changes.
Model-based reinforcement learning (RL) uses a predictive model to simulate future states and make informed decisions, enhancing the adaptability and efficiency of embedded systems. Having a model of the environment helps to adapt to new tasks with minimal data, and is particularly beneficial for embedded systems with limited computational resources and varying workloads. For example, automotive systems using multi-agent RL (MARL) can converge the DVFS strategies faster by each agent getting feedback from real-data and inference from the model of the processor, ensuring energy-efficient and reliable performance for safety-critical tasks. In mobile devices, the environment model combined with LLM-extracted semantic features enables zero-shot deployment: RL agents can be trained using synthetic data generated by the model without requiring any profiling samples from the target platform. This eliminates the sample collection overhead that limits prior transfer learning approaches. Critically, model-based approaches enable low-latency decision-making by amortizing learning costs across multiple decisions. First decisions on new benchmarks require 3.5 to 8.0s including one-time LLM feature extraction, while subsequent decisions achieve 358ms. First-decision latency is 8,300\$
times\$ faster than exhaustive table-based profiling.
@@
Through extensive experimentation and evaluation on representative embedded system platforms, we demonstrate the effectiveness of the proposed approach in achieving thermal awareness, energy efficiency, and system reliability. Several low-energy methodologies are implemented including single-agent and multi-agent direct RL and model-based RL, to be compared with federated energy-aware scheduling [7] and linux governors [9]. Our framework integrates runtime monitoring of temperature patterns using Linux in-kernel profiling and the Barcelona OpenMP Tasks Suite (BOTS) and PolybenchC for workload characterization [20]. By leveraging the capabilities of MARL, our approach offers a flexible and adaptive solution for zero-shot deployment of temperature-aware and energy-efficient scheduling on new embedded systems, paving the way for enhanced performance and sustainability in a wide range of applications such as mobile and autonomous systems.
@@
  1. 1.
    The modeling framework is integrated into the few-shot learning techniques through an off-policy MARL approach to enable efficient adaptation to new thermal scenarios with limited training data. Two separate temperature and profiler agents cooperate to increase the action space for thermal awareness and energy efficiency. A goal-based reward function based on makespan, power consumption, and temperature deviation is provided.
  2. 2.
@@
  3. 3.
    We introduce LLM-based semantic feature extraction that characterizes OpenMP programs through 13 code-level features without execution. We evaluate three state-of-the-art LLMs (DeepSeek-V3, Claude Sonnet, GPT-4o) for code analysis, achieving up to 73.8% inter-model agreement on high-level features. This enables zero-shot prediction for unseen workloads with one-time extraction cost of \$0.018 per program.
  4. 4.
    The effectiveness of the models and few-shot model-based learning approach is evaluated through extensive experiments by verified OpenMP benchmarks (BOTS and PolybenchC) and multiple embedded processors including NVIDIA Jetson TX2, Jetson Orin NX, RubikPi, and Intel Core i7. To reduce the overhead of the models to be run on a single core, we provide comprehensive tuning in terms of accuracy and architectural exploration.
  5. 5.
    The proposed scheduler is compared to the state-of-the-art heuristic-based schedulers [7] and Linux governors [9]. Our model-based approach achieves 358ms inference latency for subsequent decisions. First decisions on new benchmarks require 3.5 to 8.0s including one-time LLM feature extraction. First-decision latency is 8,300\$
@@
The paper is organized as follows: Section 2 surveys related work, Section 3 describes motivation, Section 4 presents the framework including LLM-based feature extraction, Section 5 evaluates performance including LLM feature analysis, and Section 6 concludes the paper. The appendix provides complete LLM implementation details.
@@ 2 Related Work
Few-shot learning encompasses diverse techniques like transfer learning, model-based RL, inverse RL, imitation learning, and meta-learning, all applicable to DVFS due to its sampling demands [61, 33, 64, 43, 22, 17, 23, 16, 71, 17, 5, 65, 51, 60]. While inverse RL infers reward functions from expert demonstrations [44], model-agnostic meta-learning adapts to new tasks with minimal data [21], and model-based RL approximates transition functions [42]. The works in [38, 30, 78, 76] utilize MARL, meta-state, transfer learning, and temporal dependencies on DVFS. The Q-learning algorithm in [74] turned off and on the clock on idle cycles FPGA instead of DVFS used in [19]. None of these works address few shot learning and learning overhead in RL-based DVFS governors. The existing few-shot learning approaches also need an expert demonstration or cannot directly be applied in embedded systems.
@@
Previous studies have investigated runtime sampling complexities in RL-based DVFS and proposed thermal predictive modeling for multi-core processors to predict thermal states [72, 10, 19, 25, 34, 41, 62]. The works in [38, 30] focus on temperature and utilization metrics and implement a command module on a separate neural network server. The predictive modeling in [41, 25] predicts the future thermal behavior of the multi-core processors given transient and ambient states. However, these RL training methods are computationally expensive on a single platform, and temperature models are inaccurate. This paper introduces computationally efficient and accurate modeling and a learning approach without requiring extra client-server interaction for frequency scaling.
2.1 LLM-Based Code Analysis for Performance Prediction
Recent advances in Large Language Models (LLMs) have demonstrated remarkable capabilities in understanding and reasoning about source code [15, 12]. Models such as GPT-4 [46], Claude [4], and DeepSeek-Coder [18] have been trained on extensive code corpora, enabling them to perform tasks ranging from code generation and bug detection to semantic analysis and performance optimization.
@@ 2.1 LLM-Based Code Analysis for Performance Prediction
Several studies have explored using machine learning for static code analysis and performance prediction. Laaber et al. [32] demonstrated that benchmark stability can be predicted using 58 statically-computed source code features without execution. In the HPC domain, researchers have applied neural networks to OpenMP program analysis, including predicting optimal thread counts [52] and detecting data races using graph neural networks. Recent work by Nichols et al. [45] investigated whether LLMs can predict GPU kernel performance characteristics (compute-bound vs. memory-bound) using only source code, achieving up to 64% accuracy with reasoning-capable models in zero-shot settings.
However, existing approaches for DVFS and task scheduling typically rely on runtime profiling or benchmark-specific identifiers, limiting their applicability to new workloads. Traditional static analysis tools extract syntactic features (loop counts, pragma annotations) but cannot capture semantic properties such as algorithmic complexity or memory access patterns that require understanding code behavior. This work bridges this gap by employing LLMs to extract semantic features from OpenMP programs, enabling zero-shot performance prediction for unseen workloads without the extensive profiling overhead required by existing methods.
@@ 3 Background and Motivation
3.1 Background
Here we give some background on DVFS, task-to-core allocation as optimization knobs, and the online learning based on RL for these two approaches.
@@ 3.1 Background
Deep Q-learning and Enhanced D3QN. The profiling data extracted from parallel workloads (e.g., BOTS benchmarks) is continuous, meaning the state space cannot be simply mapped to the action space through a Q-function in Q-learning. Deep Q-learning, a sophisticated RL algorithm, leverages neural networks to approximate these Q-values. Deep Q-learning provides sample efficiency by training through past experiences stored as state-action pairs inside a replay buffer. To mitigate issues like overestimation of actions, the Q-function utilizes Dueling Deep Q-Networks (D3QN), i.e., decomposing the Q-function into a value function and an advantage function, where the value function gives a global value for each corresponding state and the advantage function evaluates the advantage of taking an action in the corresponding state. The Q-value is assigned to a corresponding state-action pair by taking an action based on the same deep Q-network with respect to the target, i.e., maximized future rewards. To make the action selection independent from the module that calculates the Q-value, a parallel Deep Q-network is designed, resulting in a double dueling deep Q-Network (D3QN). These strategies collectively allow the Deep Q-Network to accurately learn optimal configurations, such as adjusting CPU frequencies and allocating tasks to specific cores. As a result, the system can dynamically manage multi-core resources, achieving improved performance and energy efficiency through a stable and data-driven policy.
@@ 3.2 Motivation on Model-Based MARL
Model-based RL provides a more efficient, adaptable, and robust framework for optimizing DVFS and task-to-core allocation in multi-core systems, achieving fast convergence and low decision latency.
3.3 Limitations of Utilization-Based DVFS
Most of the available Linux kernel governors take CPU utilization to scale voltage and frequency. Utilization is defined as:
$\textit{Utilization}=\frac{T_{\text{busy}}}{T_{\text{busy}}+T_{\text{idle}}}$
@@ 3.3 Limitations of Utilization-Based DVFS
Here, $T_{\text{busy}}$ is the time the processor is active (including both execution and stall times), and $T_{\text{idle}}$ is the idle time. The busy time $T_{\text{busy}}$ comprises both active time ($T_{\text{active}}$), when the processor is executing instructions, and stall time ($T_{\text{stall}}$), when execution is delayed due to factors like branch mispredictions, context switching overhead, cache misses, etc. [24]. Stall time can account for more than 50% of a workload’s execution time [38], yet utilization-based DVFS does not differentiate between the active and stall times. Thus, the utilization-aware frequency scaling may result in a less effective solution to match the frequency to the actual demand of the workload.
3.4 Challenges in Task-to-Core Allocation
@@
Due to the characteristics of different core types and different properties of tasks, the set and the number of cores to be allocated to the parallel workload should be carefully determined. Many modern processors support heterogeneous computing by combining cores with different characteristics to enhance power efficiency and performance capabilities. Moreover, many modern processors have non-uniform memory access (NUMA), i.e., cores have different proximity to memory locations, resulting in performance degradation for distant cores. Selecting the cores also affects thermal throttling; overloading nearby cores can stall workload execution and accelerate processor aging. Selection of the number of cores to be allocated to the parallel tasks is also important. For example, in an OpenMP workload, where each thread of execution represents a core, tied tasks should be restricted to execute on the allocated core. In contrast, untied tasks can migrate to other cores. Unpredictable execution time of a tied task, caused by branch misses, can lead to overall unpredictability in the workload’s makespan. This introduces timing unpredictability, impacting the overall makespan and energy/power consumption. As a result, increasing the number of allocated cores in this case increases unpredictability and does not lead to improved performance [3]. In sum, to make informed decisions on task-to-core allocation, observation variables such as branch misses, cache misses, context switches, and each core’s temperature should be given appropriate attention.
@@
Figure 1 highlights the changes of observation features at different frequency levels and sets of cores used for estimating the important performance metrics such as energy consumption and efficient voltage and frequency scaling and core allocation. The data is extracted from performance monitoring on an NVIDIA Jetson TX2 with 6 cores that runs the FFT benchmark parallelized through the OpenMP API. The frequency is split into 12 levels, level 0 representing the lowest frequency and level 11 representing the highest frequency. Core 0 is not allocated to the benchmark to retain system interrupt handler performance, which is mostly done at core 0. The results depict six important metrics across frequency levels and core allocation configurations. CPU utilization (Figure 1(a)) increases with more cores and shows minimal impact from frequency scaling. Makespan (Figure 1(b)) reduces with higher frequency but shows diminishing returns when cores are more than two. Energy consumption (Figure 1(c)) increases sharply with higher frequency and core count, reducing efficiency. There is a big gap in branch misses when moving from one core to multiple cores, but it is largely unaffected by frequency (Figure 1(d)). Context switches rise with more cores, indicating inefficiency in core usage (Figure 1(e)). Average temperature increases with frequency and core count, risking thermal throttling (Figure 1(f)).
Figure 1: Performance metrics (CPU utilization, makespan, energy consumption, branch misses, context switches, and average temperature) for NVIDIA Jetson TX2 running the FFT benchmark parallelized through OpenMP API. Results highlight different responses of performance metrics to changes in different frequencies and selected cores.
3.5 The Need for Environment Modeling and Multi-Agent RL
@@
To effectively address both DVFS and task-to-core allocation while enhancing sample efficiency, it is crucial to develop an accurate environment model and employ a multi-agent reinforcement learning (RL) framework. An accurate environment model captures the system’s characteristics, including hardware architecture, workload behavior, and their interactions, enabling low-cost synthetic sample generation for few-shot learning. By modeling the environment and incorporating factors such as branch misses, cache misses, and context switches, more precise and efficient strategies for energy optimization can be devised.
@@
3.6 LLM-Based Code Analysis for Workload Profiling
@@ 4.1 Design of MARL
To address this challenge, we adopt a collaborative hierarchical MARL approach, as shown in Figure 2(a). This approach decomposes the problem into lower-dimensional sub-problems handled by separate agents implemented as D3QN. The final action is a combination of each agent’s decision. We use two primary agents:
@@
(a) Hierarchical MARL Action Selection with Two Agents: The Profiler Agent determines the number of active cores and the operating frequency based on performance and energy consumption states. The Temperature Agent assigns priority levels to cores based on temperature states. The combined actions result in prioritized cores and the appropriate number allocated to tasks, reducing computational complexity and decision latency.
(b) Reward function definitions for (a) the Profiler Agent and (b) the Temperature Agent. (a) The Profiler Agent’s reward $r_{\text{profiler}}$ is based on exponential functions of energy consumption ($E_{\text{A}}$) and makespan ($\text{makespan}_{A}$). Here, a threshold factor ($c_{\text{th}}=0.3$) and a steepness factor ($c_{\text{st}}=0.5$) determine the reward focus. (b) The Temperature Agent’s reward $r_{\text{temp}}$ changes linearly based on core temperature ($\text{temp}_{i}$), aiming to maintain temperatures below the threshold of $50^{\circ}$C.
Figure 2: MARL design: (a) Hierarchical action selection architecture and (b) reward function definitions.
According to Figure 2(a), the action that assigns frequency to a core is called ($a_{freq}$) and the input energy consumption state ($s_{energy}$) determines Q-value $Q(a_{freq},s_{energy})$, while the action that assigns the number of cores is called ($a_{cores}$) based on workload performance state ($s_{workload}$) that determines Q-value $Q(a_{cores},s_{workload})$. To decrease the computational overhead, these two agents are combined resulting in profiler action $a_{profiler}:\{a_{cores},a_{freq}\}$ from the profiler state $s_{profiler}:\{s_{energy},s_{workload}\}$. Priority assignment action ($a_{temp}$) is based on the core thermal state ($s_{temp}$), determining Q-value $Q(a_{temp},s_{temp})$. Assigning a frequency from $n$ frequency levels to $m$ cores results in $m\times n$ possible actions. Selecting from $m$ possible number of cores results in $m$ choices, and for priority selection, we have choices from $m\times m$ possible actions. Introducing sequential action selection in MARL reduces this to $m\times m+m\times n+m$ possible choices, enabling inference latency under 10ms on embedded platforms.
Reward Function Definition. The reward functions for the Profiler and Temperature agents are designed to guide the system toward optimal performance and energy efficiency. Unlike previous studies [38, 30], which rely on CPU utilization and frame rate as target metrics, our approach focuses on makespan and energy consumption of suboptimal conditions defined by performance and powersave linux governors to provide a more accurate assessment of workload performance. This shift addresses the limitations of utilization metrics, which can be misleading by overlooking stall times. Additionally, to prevent thermal throttling, we enforce a temperature limit of $50^{\circ}$C based on the thermal throttling threshold of the Jetson TX2.
@@
Here, $E_{\text{total}}$ represents the total energy consumption, and makespan denotes the workload completion time. The parameters $c_{\text{th}}$ and $c_{\text{st}}$ control the threshold for penalization and the steepness of the exponential increase, respectively. By adjusting these parameters, we can prioritize either energy efficiency or makespan based on the specific optimization objectives in each agent.
Figure 2(b) illustrates these two distinct reward functions, showing how the profiler agent optimizes for energy and makespan while the temperature agent maintains thermal constraints.

### v2

# arXiv:2601.08166v2[cs.AI] 28 Feb 2026

@@ 1 Introduction
Energy-efficient embedded systems are limited by computational resources and are widely deployed in IoT devices, wearable electronics, and autonomous vehicles. In these systems, managing the thermal dynamics prevents thermal overheating, ensures system reliability, and avoids non-uniform aging due to uneven workload distribution among the cores [9, 18]. However, the available thermal-aware energy-efficient methods proposed for embedded systems lack scalability to many cores and adaptability to different platforms, especially in dynamic and resource-constrained environments, and often incur prohibitive computational overhead.
@@
Given the complexity of embedded systems, the cost-intensive nature of collecting real-world data adds to the challenges of task and processor profiling. This complexity frequently leads to unpredictable execution time estimates, aggravated by limited hardware feedback mechanisms within these systems. These challenges can significantly hinder efficient core allocation, affecting overall performance and energy efficiency. Consequently, existing task scheduling strategies mostly prioritize workload demand or only utilization metrics over temperature and system-related considerations. This trend is discernible in Linux kernel governors [9] and recently formulated policies [37, 28, 40]. As a result, these strategies often overlook detailed insights provided by system profilers during task execution [3] and allocate resource-intensive tasks to hot cores, which are more sensitive to temperature elevations. State-of-the-art heuristic approaches like the precise scheduler [7] achieve near-optimal energy efficiency for static frequency assignment through exhaustive offline profiling across all configurations, but require 8 to 12 hours of table generation per benchmark (e.g., profiling a 1 to 5 second task across all core-frequency combinations with multiple repetitions), preventing runtime adaptation. Their static nature limits adaptability to thermal variations and workload phase changes.
@@
[…]
Through extensive experimentation and evaluation on representative embedded system platforms, we demonstrate the effectiveness of the proposed approach in achieving thermal awareness, performance optimization, and system reliability. Several low-energy methodologies are implemented including single-agent and multi-agent direct RL and model-based RL, to be compared with federated energy-aware scheduling [7] and Linux governors [9]. Our framework integrates runtime monitoring of temperature patterns using Linux in-kernel profiling and the Barcelona OpenMP Tasks Suite (BOTS) and PolybenchC for workload characterization [19]. By leveraging the capabilities of MARL, our approach offers a flexible and adaptive solution for zero-shot deployment of temperature-aware and energy-efficient scheduling on new embedded systems, enabling practical deployment in mobile and autonomous systems.
@@
  • We introduce LLM-based semantic feature extraction that characterizes OpenMP programs through 13 code-level features without execution. We evaluate three state-of-the-art LLMs (DeepSeek-V3, Claude Sonnet, GPT-4o), achieving up to 73.8% inter-model agreement, enabling zero-shot deployment to new platforms with one-time extraction cost of \$0.018 per program.
  • We develop thermal-aware accurate modeling that captures thermal dynamics of embedded processors under varying workloads. The proposed regression models outperform state-of-the-art temperature prediction by a factor of over 6 in Mean Squared Error (MSE), with model inference latencies under 5ms.
@@ 2 Related Work
Few-shot learning encompasses diverse techniques like transfer learning, model-based RL, inverse RL, imitation learning, and meta-learning, all applicable to DVFS due to its sampling demands [64, 31, 67, 42, 21, 16, 5, 68, 54]. While inverse RL infers reward functions from expert demonstrations [43], model-agnostic meta-learning adapts to new tasks with minimal data [20], and model-based RL approximates transition functions [41]. The works in [37, 28, 80, 78] utilize MARL, meta-state, transfer learning, and temporal dependencies on DVFS. The Q-learning algorithm in [76] performs clock gating on idle FPGA cycles rather than using DVFS as in [18]. None of these works address few shot learning and learning overhead in RL-based DVFS governors. The existing few-shot learning approaches also need an expert demonstration or cannot directly be applied in embedded systems.
@@
Previous studies have investigated runtime sampling complexities in RL-based DVFS and proposed thermal predictive modeling for multi-core processors to predict thermal states [74, 10, 18, 23, 32, 40, 65]. The works in [37, 28] focus on temperature and utilization metrics and implement a command module on a separate neural network server. The predictive modeling in [40, 23] predicts the future thermal behavior of the multi-core processors given transient and ambient states. However, these RL training methods are computationally expensive on a single platform, and temperature models are inaccurate. Building upon hierarchical multi-agent DVFS [50], feature-aware statistical learning [51], and graph-driven performance modeling [49], this paper introduces computationally efficient and accurate modeling combined with flow-augmented few-shot learning techniques [52] and LLM-based semantic feature extraction, enabling a learning approach without requiring extra client-server interaction for frequency scaling.
@@
Recent advances in Large Language Models (LLMs) have demonstrated remarkable capabilities in code understanding and performance prediction [15, 12, 45, 4, 17]. Laaber et al. [30] demonstrated that benchmark stability can be predicted using 58 statically-computed source code features without execution, while Sanz et al. [55] applied neural networks to predict optimal thread counts for OpenMP regions. Recent work by Nichols et al. [44] investigated whether LLMs can predict GPU kernel performance characteristics using only source code, achieving up to 64% accuracy with reasoning-capable models in zero-shot settings. However, existing approaches for DVFS and task scheduling typically rely on runtime profiling or benchmark-specific identifiers, limiting their applicability to new platforms and workloads. This work bridges this gap by employing LLMs to extract semantic features from OpenMP programs that, combined with environment models, enable zero-shot deployment to new platforms without platform-specific profiling overhead.
@@ 3 Background and Motivation
Deep Q-learning and Enhanced D3QN. The profiling data extracted from parallel workloads (e.g., BOTS benchmarks) is continuous, meaning the state space cannot be simply mapped to the action space through a Q-function in Q-learning. Deep Q-learning leverages neural networks to approximate these Q-values. Deep Q-learning provides sample efficiency by training through past experiences stored as state-action pairs inside a replay buffer. To mitigate issues like overestimation of actions, the Q-function utilizes Dueling Deep Q-Networks (D3QN), i.e., decomposing the Q-function into a value function and an advantage function, where the value function gives a global value for each corresponding state and the advantage function evaluates the advantage of taking an action in the corresponding state. The Q-value is assigned to a corresponding state-action pair by taking an action based on the same deep Q-network with respect to the target, i.e., maximized future rewards. To make the action selection independent from the module that calculates the Q-value, a parallel Deep Q-network is designed, resulting in a double dueling deep Q-Network (D3QN). These strategies allow the Deep Q-Network to learn optimal configurations, such as adjusting CPU frequencies and allocating tasks to specific cores. As a result, the system can dynamically manage multi-core resources, achieving improved performance and energy efficiency through a stable and data-driven policy.
@@
Motivation on Model-Based MARL. Model-based RL provides a more efficient, adaptable, and robust framework for optimizing DVFS and task-to-core allocation in multi-core systems, achieving fast convergence and low decision latency.
@@
Limitations of Utilization-Based DVFS. Most of the available Linux kernel governors take CPU utilization to scale voltage and frequency. Utilization is defined as $\textit{Utilization}=T_{\text{busy}}/(T_{\text{busy}}+T_{\text{idle}})$, where $T_{\text{busy}}$ is the time the processor is active (including both execution and stall times), and $T_{\text{idle}}$ is the idle time. The busy time $T_{\text{busy}}$ comprises both active time ($T_{\text{active}}$), when the processor is executing instructions, and stall time ($T_{\text{stall}}$), when execution is delayed due to factors like branch mispredictions, context switching overhead, cache misses, etc. [22]. Stall time can account for more than 50% of a workload’s execution time [37], yet utilization-based DVFS does not differentiate between the active and stall times. Thus, the utilization-aware frequency scaling may result in a less effective solution to match the frequency to the actual demand of the workload.
Figure 1: Performance metrics for FFT on Jetson TX2 across frequency/core configurations.
@@
Challenges in Task-to-Core Allocation. Due to the characteristics of different core types and different properties of tasks, the set and the number of cores to be allocated to the parallel workload should be carefully determined. Many modern processors support heterogeneous computing by combining cores with different characteristics to enhance power efficiency and performance capabilities. Moreover, many modern processors have non-uniform memory access (NUMA), i.e., cores have different proximity to memory locations, resulting in performance degradation for distant cores. Selecting the cores also affects thermal throttling; overloading nearby cores can stall workload execution and accelerate processor aging. Selection of the number of cores to be allocated to the parallel tasks is also important. For example, in an OpenMP workload, where each thread of execution represents a core, tied tasks should be restricted to execute on the allocated core. In contrast, untied tasks can migrate to other cores. Unpredictable execution time of a tied task, caused by branch misses, can lead to overall unpredictability in the workload’s makespan. This introduces timing unpredictability, impacting the overall makespan and energy/power consumption. As a result, increasing the number of allocated cores in this case increases unpredictability and does not lead to improved performance [3]. In sum, to make informed decisions on task-to-core allocation, observation variables such as branch misses, cache misses, context switches, and each core’s temperature should be given appropriate attention.
@@
Figure 1 highlights observation features at different frequency levels and core counts on NVIDIA Jetson TX2 running FFT. The frequency has 12 levels (0=lowest, 11=highest). CPU utilization increases with cores but shows minimal frequency impact. Makespan reduces with higher frequency but shows diminishing returns beyond two cores. Energy consumption increases sharply with frequency and core count. Branch misses exhibit a large gap between single and multiple cores but remain frequency-independent. Context switches rise with cores, indicating allocation inefficiency. Temperature increases with frequency and core count, risking thermal throttling.
Environment Modeling and Multi-Agent RL.
@@
To effectively address both DVFS and task-to-core allocation while enhancing sample efficiency, we develop an accurate environment model and employ a multi-agent reinforcement learning (RL) framework. An accurate environment model captures the system’s characteristics, including hardware architecture, workload behavior, and their interactions, enabling low-cost synthetic sample generation for few-shot learning. By modeling the environment and incorporating factors such as branch misses, cache misses, and context switches, more precise and efficient strategies for energy optimization can be devised.
@@
LLM-Based Workload Profiling. Traditional workload characterization relies on either exhaustive runtime profiling (8-12 hours per platform) or syntactic static analysis that captures code structure but not semantic meaning. Recent Large Language Models (LLMs) such as GPT-4 [45], Claude [4], and DeepSeek [17] enable zero-shot semantic feature extraction from source code without execution. LLMs can infer algorithmic complexity, memory access patterns, parallelization characteristics, and synchronization overhead by analyzing code semantics. This enables performance prediction for unseen workloads and cross-platform transfer learning, as semantic features (e.g., $O(n^{2})$ complexity, unit-stride access) remain platform-agnostic. We extract 13 semantic features across memory, algorithmic, and parallelization dimensions to complement hardware performance counters, achieving MAPE under 15% for zero-shot predictions.
@@ 4.1 Design of MARL
To address this challenge, we adopt a collaborative hierarchical MARL approach that decomposes the exponential action space into lower-dimensional sub-problems handled by separate agents implemented as Double Deep Q-Networks (D3QN). The final action is a combination of each agent’s decision. We use two primary agents:
@@
In the hierarchical decomposition, the action that assigns frequency to a core is called ($a_{freq}$) and the input energy consumption state ($s_{energy}$) determines Q-value $Q(a_{freq},s_{energy})$, while the action that assigns the number of cores is called ($a_{cores}$) based on workload performance state ($s_{workload}$) that determines Q-value $Q(a_{cores},s_{workload})$. To decrease the computational overhead, these two agents are combined resulting in profiler action $a_{profiler}:\{a_{cores},a_{freq}\}$ from the profiler state $s_{profiler}:\{s_{energy},s_{workload}\}$. Priority assignment action ($a_{temp}$) is based on the core thermal state ($s_{temp}$), determining Q-value $Q(a_{temp},s_{temp})$. Assigning a frequency from $n$ frequency levels to $m$ cores results in $m\times n$ possible actions. Selecting from $m$ possible number of cores results in $m$ choices, and for priority selection, we have choices from $m\times m$ possible actions. Introducing sequential action selection in MARL reduces this to $m\times m+m\times n+m$ possible choices, enabling inference latency under 10ms on embedded platforms.
Reward Function Definition. The reward functions for the Profiler and Temperature agents are designed to guide the system toward optimal performance and energy efficiency. Unlike previous studies [37, 28], which rely on CPU utilization and frame rate as target metrics, our approach focuses on makespan and energy consumption of suboptimal conditions defined by performance and powersave linux governors to provide a more accurate assessment of workload performance. This shift addresses the limitations of utilization metrics (defined as $\text{utilization}=t_{\text{busy}}/(t_{\text{busy}}+t_{\text{idle}})$), which can be misleading by overlooking stall times. Additionally, to prevent thermal throttling, we enforce a temperature limit of $50^{\circ}$C based on the thermal throttling threshold of the Jetson TX2.
@@
Here, $E_{\text{total}}$ represents the total energy consumption, and makespan denotes the workload completion time. The parameters $c_{\text{th}}$ and $c_{\text{st}}$ control the threshold for penalization and the steepness of the exponential increase, respectively. By adjusting these parameters, we can prioritize either energy efficiency or makespan based on the specific optimization objectives in each agent. These distinct reward functions guide the profiler agent to optimize for energy and makespan while the temperature agent maintains thermal constraints.